Financial data for Joel de Paris, Inc., for last year follow:

Joel de Paris, Inc.
Balance Sheet
Beginning
Balance
Ending
Balance
Assets
Cash $ 128,000 $ 136,000
Accounts receivable 339,000 478,000
Inventory 561,000 483,000
Plant and equipment, net 892,000 863,000
Investment in Buisson, S.A. 405,000 431,000
Land (undeveloped) 251,000 245,000
Total assets $ 2,576,000 $ 2,636,000
Liabilities and Stockholders' Equity
Accounts payable $ 371,000 $ 334,000
Long-term debt 983,000 983,000
Stockholders' equity 1,222,000 1,319,000
Total liabilities and stockholders' equity $ 2,576,000 $ 2,636,000


Joel de Paris, Inc.
Income Statement
Sales $ 4,462,000
Operating expenses 3,881,940
Net operating income 580,060
Interest and taxes:
Interest expense $ 113,000
Tax expense 206,000 319,000
Net income $ 261,060


The company paid dividends of $164,060 last year. The “Investment in Buisson, S.A.,” on the balance sheet represents an investment in the stock of another company. The company's minimum required rate of return of 15%.

Required:

1. Compute the company's average operating assets for last year.

2. Compute the company’s margin, turnover, and return on investment (ROI) for last year. (Do not round intermediate calculations and round your final answers to 2 decimal places.)

3. What was the company’s residual income last year?
Answer #1

Average operating assets

= [(2576000-251000-405000)+(2636000-245000-431,000)]/2

= 1,940,000

Margin = Net operating income/Sales

= 580060/4462000

= 13%

Turnover = Sales/Average operating assets

= 4462000/1940000

= 2.3

Roi = Margin*Turnover

= 13*2.3

= 29.9%

Residual income = Net operating income - Required return

= 580,060-(1,940,000*15%)

= 289,060